OpenStudy (anonymous):

Initial value problem, dy/dt + y/t = g(t), y(1)=1 No idea where to start.

OpenStudy (anonymous):

Don't you find the general solution and then find the particular solution?

OpenStudy (anonymous):

Actually I noticed something interesting...

OpenStudy (anonymous):

I don't know how to begin getting the general solution, that's the problem. :( All the questions we've done so far have had only one function in them. For instance, I just solved \[dy/dt=e^{t-y}\] but that only has y as a function of t. This new question has both g AND y as functions of t.

OpenStudy (anonymous):

\[ \begin{array}{rcl} \frac{dy}{dt} + \frac{y}{t} &=& g(t) \\ \frac{dy}{dt}\cdot t + y\cdot 1 &=& t\cdot g(t) \\ \frac{d}{dt}(y\cdot 1)&=& t\cdot g(t) \end{array} \]

OpenStudy (anonymous):

oops should be \[ \frac{d}{dt}\left( y\cdot t\right) =g(t)\cdot t \]

OpenStudy (anonymous):

By the way, do you know how to do something like the Laplace transform?

OpenStudy (anonymous):

Anyway, when getting the general solution, It should be separable.

OpenStudy (anonymous):

\[ \begin{array}{rcl} \frac{dy}{dt} + \frac{y}{t} &=& 0 \\ \frac{dy}{dt} &=& - \frac{y}{t} \\ \frac{1}{y}\frac{dy}{dt} &=& - \frac{1}{t} \\ \int \frac{1}{y}\frac{dy}{dt} dt &=& \int - \frac{1}{t} dt \\ \int \frac{1}{y} dy &=& \int - \frac{1}{t} dt \\ \end{array} \]

OpenStudy (anonymous):

Never heard of the Laplace transformation, but that's sort of why I'm studying right now, haha

OpenStudy (anonymous):

Anyway, see how the general solution is just \(\frac{C}{t}\)